Find the slope of a line containing these points.
(5,-2) & (-3,2)
need help and steps

Answer by Alan3354(69443)

Find the slope of a line containing these points.
(5,-2) & (-3,2)
----------
Always work left to right, so
Point 1 is (-3,2)
Point 2 is (5,-2)
Find the slope, m
m = (y2-y1)/x2-x1)
m = (-2-2)/(5-(-3))
m = -4/8
m = -1/2 (the slope)
It's important to "work left to right", from small value of x to larger. This is a convention, meaning it's done by agreement so we all mean the same thing when we calculate slope.
-------------
To find the eqn:
Use either point to find the equation, I'll use Point 1
y-y1 = m(x-x1)
y-2 = (-1/2)*(x-(-3))
y-2 = -(x/2) - 3/2
2y-4 = -x - 3
x+2y = 1